How many grams of h20 are produced from 0.25 moles of o2


1
Expert's answer
2021-02-25T03:50:27-0500

2H2 + O2 = 2H2O


n(O2):n(H2O)=1:2

n(H2O)=0.5mol


m(H2O)=0.5×18=9g